If f(x)=2x-1 and g(x)=x-2, find f(g(x)) and g(f(x).
Answer by sachi(548)

If f(x)=2x-1 and g(x)=x-2, find f(g(x)) and g(f(x).

ans:-
f(g)x))=f(x-2)=2(x-2)-1=2x-4-1=2x-5
g(f(x))=g(2x-1)=2x-1-2=2x-3
ans